(a) A 2.00-uF capacitor is connected to a 9.00-V battery. How much energy is stored in the capacitor? (b) Had the capacitor been connected to a 6.00-V battery, how much energy would have been stored?
Answer #1

a] Energy = 0.5 CV^2 = 0.5*2*9^2 = 81 uJ answer

b] Energy = 0.5 CV^2 = 0.5*2*6^2 = 36 uJ answer